The area of lawn is 1850m^2. Its length 6 time from its width. Its width will be
Answer by checkley77(12844)

AREA=L*W
L=6W
1850=6W*W
1,850=6W^2
W^2=1,850/6
W^2=308.33
W=SQRT308.33
W=17.56 METERS. FOR THE WIDTH.
PROOF;
1,850=6*17.56*17.56
1,850=1,850